OpenStudy (anonymous):

How to rewrite the this as f(x)=mx+b y+x+8=0 i wish i can get full procedure

hartnn (hartnn):

so you need to isolate y from y+x+8=0 just subtract x+8 from both sides, and you will get 'y' on left side isolated.

OpenStudy (anonymous):

can i see how should i do it?

hartnn (hartnn):

let me give you an example, then you try this one :) 2x+y -12 =0 now i have to isolate 'y' , so i will subtract (2x-12) from both sides, \(\Large 2x+y-12 -(2x-12) = 0-(2x-12) \\ \Large 2x-2x +y -12+12 = -2x+12 \\ \Large 0+y +0 =-2x+12 \\ \Large y = -2x+12\) do the same procedure for y+x+8 =0 try it, if you make an error, we are her :)

hartnn (hartnn):

here** lol

OpenStudy (anonymous):

y+x+8=0 (-x+8) +y+x+8= 0-x+8 y=0-x+8 y=-x+8 how about this? correct me if im wrong please thnx xD

hartnn (hartnn):

just one error : you subtracted -x+8 you should have subtracted -(x+8) or -x-8

OpenStudy (anonymous):

how should i write it please help me

hartnn (hartnn):

y+x+8=0 (-x-8) +y+x+8= 0-x-8 y=0-x-8 y=-x-8 thats it!

OpenStudy (anonymous):

@shiqki017 are you aware of transposing terms? Like bringing terms from L.H.S to R.H.S?

OpenStudy (anonymous):

what do you mean sir?

OpenStudy (anonymous):

y+x+8=0 So as @hartnn said.. you have to isolate 'y'.... So transpose r to the rhs! RHS is the part to the right of the equation that is 0 in this case. LHS is the part to the left of the equation that is (y+x+8) in this case. While transposing or moving terms from the RHS to LHS or from the LHS to RHS, positives become negative and vice-versa. So, here in this equation y+x+8=0 transpose x+8 from the LHS to RHS to isolate y on the LHS. So, the equation becomes y=-x-8+0 which is y=-x-8. If this is confusing, follow @hartnn's method. :)

OpenStudy (anonymous):

@shiqki017 there is a typo in the third line *So transpose x+8 to the RHS. Sorry.